3. Gains from trade Consider two neighboring island countries called Dolorium and Contente. They each have 4 million labor hours available per week that they can use to produce jeans, rye, or a combination of both. The following table shows the amount of jeans or rye that can be produced using 1 hour of labor. Jeans Rye (Bushels per hour of labor) 20 16 (Pairs per hour of labor) Country Dolorium Contente Initially, suppose Contente uses 1 million hours of labor per week to produce jeans and 3 million hours per week to produce rye, while Dolorium uses 3 million hours of labor per week to produce jeans and 1 million hours per week to produce rye. Consequently, Dolorium produces 15 million pairs of jeans and 20 n bushels of rye, and Contente produces 8 million pairs of jeans and 48 million bushels of rye. Assume there are no other countries willing to trade goods, so, in the absence of trade between these two countries, each country consumes the amount of jeans and rye it produces. Dooriums opportunity cost of producing 1 pair of eans is 4 bushels ▼ of rye, and Contentes opportunity cost of producing 1 pair of jeans is 2 bushels , of rye. Therefore, D lori m ▼ has a comparative advantage in the production of eans, and Contente has a comparative advantage in the production of rye. Suppose that each country completely specializes in the production of the good in which it has a comparative advantage, producing only that good. In this case, the country that produces jeans will produ ce11 million pairs per week, and the country that produces rye will produce 52 million bushels per week. In the following table, enter each countrys production decision on the third row of the table (marked Production) Suppose the country that produces jeans trades 18 million pairs of jeans to the other country in exchange for 54 million bushels of rye.Homework (Ch 03) In the foilowing table, enter each countrys production decision on the third row of the table (manked Production). Suppose the country that produces jeans trades 18 million pairs of jeans to the other country in exchange for 54 million bushels of rye. In the following table, select the amount of each good that each country exports and imports in the boxes across the row marked Trade Action, and enter each countrys final consumption of each good on the line marked Consumption. When the two countries did not specialize, the total production of jeans was 23 million pairs per week, and the total production of rye was 68 million bushels per week. Because of specialization, the total production of jeans has increased by rye has increased by million pairs per week, and the total production of million bushels per week. Because the two countries produce more jeans and more rye under specialization, each country is able to gain from trade Calculate the gains from trade-that is, the amount by which each country has increased its consumption of each good relative to the first row of the table. In the following table, enter this difference in the boxes across the last row (marked Increase in Consumption). Dolorium Contente Rye Jeans Rye (Millions of pairs) (Millions of bushels) (Millions of pairs) (Millions of bushels) Without Trade Production 15 20 48 Consumption 15 20 48 With Trade Production Trade action Consumption Gains from Trade Increase in Consumption
Answer #1

Dolorium:

5 pairs of Jeans = 20 bushels of Rye

1 pair of Jeans = 20/5 = 4 bushels of Rye

Opportunity cost of producing 1 pair of Jeans is 4 bushels of Rye.

Contente:

8 pairs of Jeans = 16 bushels of Rye

1 pair of Jeans = 16/8 = 2 bushels of Rye

Opportunity cost of producing 1 pair of Jeans is 2 bushels of Rye.

A country has a comparative advantage in producing that good if the opportunity cost of producing that good is lower in that country as compared to another country.

Opportunity cost of producing 1 pair of Jeans is lower in Contente so, Contente has a comparative advantage in Jeans and Dolorium has a comparative advantage in Rye.

When Contente uses 4 million hours of labor to produce Jeans then it is able to produce 4 million X 8 = 32 million pairs of jeans

When Dolorium uses 4 million hours of labor to produce Rye then it is able to produce 4 million X 20 = 80 million bushels of Rye

Total production of Rye increased by 80 million - (20 + 48) million = 12 million

Total production of Jeans increased by 32 million - (15 + 8) million = 9 million

Dolorium Contente Jeans Rye Jeans Rye (Millions of pairs) (Millions of bushels) (Millions of pairs) (Millions of bushels) Wit